# Cascade-Zero123: One Image to Highly Consistent 3D with Self-Prompted Nearby Views

### v1

@@ A.2 More Qualitative Results of Novel View Synthesis
We show more qualitative results generated by our Cascade-Zero123 in Fig. 6. Our model can generalize well to unseen data. We selected more various images from real-world scenes or high-quality scenes generated by Stable Diffusion 2.1 [46] for the experiments on novel view synthesis. These scenes include different kinds of environments and objects. We tested more scenarios involving multiple object stacking (such as two geese, a computer, and a pumpkin) and complex branching structures (such as a satellite). The selected example images were also deliberately chosen to avoid central symmetry or left-right symmetry, which can pose challenges for generating novel views using conditional latent diffusion models. In the case of these difficult-to-maintain consistency scenes, our Cascade-Zero123 achieved better consistency compared to Zero123-XL and significantly improved image quality compared to SyncDreamer [31]. For SyncDreamer, we choose a variety of elevation angles and random seeds to generate images that try our best to show better performance.
A.3 Cost Analysis
@@
As shown in Table. 4, we compared the pre-training time per iteration between Zero-1-to-3 and our Cascade-Zero123, on the same machine and environment. It can be observed that our model does not significantly increase the pre-training time. Additionally, during the Single image-to-3D stage, there is only a slight additional inference cost of around a couple of seconds for Base-0123. This extra time can be negligible for single image-to-3D tasks that typically take dozens of minutes.
@@
Methods $\downarrow$ Pretraining Time
Zero-1-to-3 [29] 9 seconds
Ours 15.7 seconds
A.4 The Pose Transition of the Objaverse Dataset Coordinate
The pose transition of the Objaverse dataset [10] coordinate is set as follows:
$\left\{\begin{matrix}x,y,z&=&R^{T}\times T\\ \theta&=&arctan(\frac{y}{x})\\ \phi&=&arctan(\frac{z}{\sqrt{(x^{2}+y^{2})}})\\ r&=&\sqrt{(x^{2}+y^{2}+z^{2})}\\ \end{matrix}\right.$ (14)
where the spherical coordinate $\theta$ is the azimuth angle measured in radians from the positive x-axis to the projection, $\phi$ is the elevation angle defined from the Z-axis down and r is the length of the line from the origin to the camera point). Then $(R^{t},T^{t})\ominus(R^{c},T^{c})$ can be formulated in the spherical coordinate as $(\theta_{t}-\theta_{c},\phi_{t}-\phi_{c},r_{t}-r_{c})$.
A.5 Discussion
We will also discuss the possibility of the cascade of view-conditional latent diffusion models in more detail. Models that prioritize consistency and higher quality than Zero-1-to-3 can also benefit from our cascade framework by constructing prompts from multiple views. In our future works, we will experiment with, for example, Cascade SyncDreamer, Cascade Magic123, and Cascade Wonder3D which can also benefit from generating prompts from additional views and cascade structure. Additionally, different types of models can also be cascaded, such as cascading a Consistent123 with Zero123.

### v2

@@ 0.A.3 More Qualitative Results of Novel View Synthesis
We show more qualitative results generated by our Cascade-Zero123 in Fig. 7. Our model can generalize well to unseen data. We selected more various images from real-world scenes or high-quality scenes generated by Stable Diffusion 2.1 [65] for the experiments on novel view synthesis. These scenes include different kinds of environments and objects. We tested more scenarios involving multiple object stacking (such as two geese, a computer and a pumpkin) and complex branching structures (such as a satellite). The selected example images were also deliberately chosen to avoid central symmetry or left-right symmetry, which can pose challenges for generating novel views using conditional latent diffusion models. In the case of these difficult-to-maintain consistency scenes, our Cascade-Zero123 achieved better consistency compared to Zero123-XL and significantly improved image quality compared to SyncDreamer [46]. For SyncDreamer, we choose a variety of elevation angles and random seeds to generate images that try our best to show better performance.
0.A.4 Cost Analysis
@@
As shown in Table. 8, we compared the pre-training time per iteration between Zero-1-to-3 and our Cascade-Zero123, on the same machine and environment. It can be observed that our model does not significantly increase the pre-training time. As for inference costs, generating one view costs 22 seconds for Zero-1-to-3 and 101 seconds for ours. However, generating 100 views as an example, Zero-1-to-3 takes 2203 seconds while our method takes 2356 seconds (6.9% increased) in total, as generating self-prompted views from the first stage is once-for-all which can be reused. Additionally, during the Single image-to-3D stage, there is also only a slight additional inference cost.
This extra time can be negligible for single image-to-3D tasks that typically take dozens of minutes.
@@
[…]
Zero-1-to-3 [43] 9 seconds 22 seconds 2203 seconds
Ours 15.7 seconds 101 seconds 2356 seconds
0.A.5 Discussion
The advantage and contribution of Cascade-Zero123 can be summarized that it can progressively dig out the multi-view information but in a reliable range. Cascade design is effective and is also adopted in other diffusion models, e.g. SDXL [59]. More reliable views can lead to better performance in reconstruction models as in PixelNeRF [106]. We also prove this in generation models via the cascade design by applying the multi-view information on the DM condition. Compared with directly using Zero 1-to-3 twice, the cascade strategy limits the prompting view in a reliable range. Our ablation study in Table LABEL:tab:ablation also demonstrates that using unreliable views as conditions can lead to negative effects. It may introduce errors and inconsistencies in the generation process. Also, fine-tuning the cascaded model equips it with a capacity to handle multi-view conditions while the original Zero 1-to-3 can only take a single view as the condition. Furthermore, compared with One-2-3-45 [41], our designed cascade method shows superior performance.